# Learning Encoding-Decoding Direction Pairs to Unveil Concepts of Influence in Deep Vision Networks

### v1

@@ 4.3 Signal Vectors as Concept Signal Estimators
In this paragraph, we consider the new data model that was outlined in Section 4.1 and try to validate whether (1) can be used to estimate the signal direction of a concept. Suppose that our objective is to estimate the signal direction of concept $i$ when we have access to a collection of patch embeddings $\{{\bm{x}}_{{\bm{p}}}\}$ and their signal values $\{\alpha_{{\bm{p}},i}\}$. Starting from the approach of Kindermans et al. (2017), it is easy to prove that whenever ${a}_{{\bm{p}},i}$ is independent of all ${a}_{{\bm{p}},j},j\neq i$ and ${\beta}_{{\bm{p}},f}$ the following property holds: $\text{cov}[{\bm{x}}_{{\bm{p}}}-{a}_{{\bm{p}},i}{\bm{s}}_{i},{a}_{{\bm{p}},i}]=0$, which, according to Kindermans et al. (2017) directly implies that (1) can be utilized to estimate the direction of the signal.
@@ 4.4 Uncertainty Region Alignment to Discover Meaningful Concepts of Influence
To accomplish the alignment, all the patch embeddings ${\bm{x}}_{{\bm{p}}}$ in an image are manipulated towards the direction $\bm{dx_{{\bm{p}}}}$ to arrive at ${\bm{x}}_{{\bm{p}}}^{\prime}={\bm{x}}_{{\bm{p}}}-\bm{dx_{{\bm{p}}}}$. Based on our estimates of ${\bm{w}}_{i}$, $b_{i}$, and $\hat{{\bm{s}}}_{i}$, we select the direction $\bm{dx_{{\bm{p}}}}$ so that the shifted ${\bm{x}}_{{\bm{p}}}^{\prime}$ lies at the intersection of the concept detectors’ decision hyperplanes. Then, we ensure the network’s prediction for the resulting manipulated image representation is highly uncertain, effectively aligning both uncertainty regions. More specifically, we define two types of feature manipulation for this purpose:
Unconstrained Feature Manipulation (UFM) in which ${\bm{x}}_{{\bm{p}}}^{\prime}=g_{{\bm{p}}}({\bm{x}}_{{\bm{p}}};{\bm{W}})={\bm{x}}_{{\bm{p}}}-\bm{dx_{{\bm{p}}}}$ and $\bm{dx_{{\bm{p}}}}$ such that:
with $\bm{A}$$^{+}$ denoting the pseudo-inverse of $\bm{A}$}.
Constrained Feature Manipulation (CFM) in which we avoid manipulating features towards datapoints that fall outside the concept encoding manifold of the network, by restricting feature manipulation to occur within the span of the signal vectors, i.e. $\bm{dx_{{\bm{p}}}}=\hat{{\bm{S}}}{\bm{v}}_{{\bm{p}}}$, ${\bm{v}}_{{\bm{p}}}\in{\mathbb{R}}^{I}$, with $\hat{{\bm{S}}}\in{\mathbb{R}}^{D\times I}$ denoting the matrix whose columns correspond to the learned signal vectors $\hat{{\bm{s}}}_{i},i\in\{0,1,...,I-1\}$. The feature manipulation formula for a patch embedding ${\bm{x}}_{{\bm{p}}}$ is $h_{{\bm{p}}}({\bm{x}}_{{\bm{p}}};{\bm{W}},\hat{{\bm{S}}})={\bm{x}}_{{\bm{p}}}-\bm{dx_{{\bm{p}}}}$, with $\bm{dx_{{\bm{p}}}}$ given by:
@@
with $\mathcal{H}$} denoting entropy and $f^{+}$ denoting the part of the network after the layer of study that provides output class probabilities. When we manipulate the representations via either UFM or CFM, we use the layer’s activation function (typically a ReLU) to keep the features in the input domain of the next layer. Throughout the experiments, we will denote Uncertainty Region Alignment loss with UFM as $\mathcal{L}^{uur}$ and with CFM as $\mathcal{L}^{cur}$. We consider the use of $\mathcal{L}^{uur}$ and $\mathcal{L}^{cur}$ to be mutually-exclusive. We either use the first or the second type of manipulation when computing the Uncertainty Region Alignment Loss and never both types at the same time.
@@ 4.5 Augmented Lagrangian Loss for Effective Direction Learning
Most of the loss terms that we use in our method are highly antagonistic. For example, concept detectors can make more confident predictions ($\mathcal{L}^{ma}$) when the separation margin ($\mathcal{L}^{mm}$) between positive and negative samples is small (while we seek to maximize it). In our experiments, we also observed that a high separation margin is also antagonistic to the sparsity of the clustering ($\mathcal{L}^{s}$ or $\mathcal{L}^{fs}$). Therefore, if we linearly combine the loss terms, the implementation of our approach across various network architectures and datasets necessitates meticulous adjustment of the loss weights within the objective function. The latter can be particularly laborious and time-consuming. Additionally, adjusting the loss weights only indirectly controls the quality of the clustering, as the final loss values can vary significantly with different weight sets. For these reasons, we choose a different strategy to learn our direction pairs. We base our method on $\epsilon$-constrained optimization, formulating our objective as a minimization problem with inequality constraints. We use the Augmented Lagrangian formulation (Hestenes, 1969; Bertsekas, 2014) to solve the following optimization problem:
$\min\quad\lambda^{fs}\mathcal{L}^{fs}+\lambda^{ur}\mathcal{L}^{ur}\quad s.t.\quad\begin{cases}\mathcal{L}^{ma}\leq\tau^{ma}\\ \mathcal{L}^{ic}\leq\tau^{ic}\\ \mathcal{L}^{mm}\leq\tau^{mm}\\ \mathcal{L}^{eac}\leq\tau^{eac}\\ \mathcal{L}^{fso}\leq\tau^{fso}\\ \mathcal{L}^{cs}\leq\tau^{cs}\end{cases}$ (13)
@@ 5.1 Reading Concept Information and Intervening on their Encoding
Reading Concept Information from the Embeddings Reading concept information from a patch embedding ends up in estimating the concept’s signal value that is encoded in the representation. Supposing that the related conditions are met, in Section 3.2 it was shown that the filter direction can serve as a means to extract this value. While in the data-model of Section 3.2, the extraction of the signal’s value is exact, under the multi-concept signal-distractor data model of Section 4.1 the considered latent space bias $\bm{c}$} is interfering with this value. More specifically, if we suppose ${\bm{w}}_{i}:{\bm{w}}_{i}\perp\hat{{\bm{s}}}_{j},i\neq j,{\bm{w}}_{i}^{T}\hat{{\bm{s}}}_{i}=1$ and ${\bm{w}}_{i}\perp{\bm{D}}$, the inner product between the filter and the embedding results in:
Thus, when considering our extended data model, the filter can be used to extract the signal’s value but with an offset that is constant, regardless of ${\bm{x}}_{{\bm{p}}}$. Due to the offset being constant, the difference between two projected embeddings is equal to the difference of their signal values, e.g.:
And second, the concept’s point of uncertainty. Given a query point ${\bm{x}}_{{\bm{p}}}$, the concept’s point of uncertainty $\hat{{\bm{u}}}_{{\bm{p}}}$ is the point that lies on the concept detector’s hyperplane (thus corresponds to a point of maximum concept ambiguity) and the line that is defined by the query point and the direction of the signal vector. More specifically, the point of uncertainty is equal to $\hat{{\bm{u}}}_{{\bm{p}}}={\bm{x}}_{{\bm{p}}}-\kappa_{{\bm{p}},i}\hat{{\bm{s}}}_{i}$ with $\kappa_{{\bm{p}},i}\in{\mathbb{R}}$ corresponding to the solution of the following system of equations:
Since (18) corresponds to the function of the concept detector, this difference is positive whenever the concept is present in ${\bm{x}}_{{\bm{p}}}$ and negative whenever the concept is absent. Thus, centering the signal value of a patch embedding around its point of uncertainty can be simply done by computing $\kappa_{{\bm{p}},i}$ using (17).
Intervening on Concept Encoding Altering the concept content that is encoded in a patch embedding ends up overwriting the signal value of the concept with a target value of interest. The target value is meant to be copied from a target embedding. Let ${\bm{x}}_{{\bm{p}}}$ denote the patch embedding that we aim to intervene on and ${\bm{x}}_{{\bm{p}}}^{t}$ denote the target embedding. Let ${\bm{x}}_{{\bm{p}}}^{\prime}$ denote the altered embedding after the intervention. Supposing that we aim to intervene on the value of concept $i$, ${\bm{x}}_{{\bm{p}}}^{\prime}$ becomes:
${\bm{x}}_{{\bm{p}}}^{\prime}={\bm{x}}_{{\bm{p}}}+\kappa_{{\bm{p}},i}\hat{{\bm{s}}}_{i}$ (19)
@@
5.2.1 Terminology and Definitions
Suppose that we want to explain the prediction of an image classifier $f$ for image $\mathcal{I}$. Let $\bm{X}$} denote this image’s representation at the penultimate layer of the network (in CNNs, this is typically the last convolutional layer before GAP). Let also ${\bm{x}}_{{\bm{p}}}\in{\mathbb{R}}^{D}$ denote a patch embedding in $\bm{X}$} at the spatial location ${\bm{p}}=(w,h)$. Following the previously introduced methodology, we assume that we have learned Encoding-Decoding Direction Pairs for this layer, i.e. filters $\bm{W}$} and signal vectors $\hat{{\bm{S}}}$. Let ${l}_{c}\in{\mathbb{R}}$ denote the network’s output logit for class $c$ regarding input $\mathcal{I}$. Typically, $c$ will be chosen such that ${l}_{c}$ is the maximum logit among the output class logits, but other choices are also valid, for instance when we want to explain why the prediction deviated from a specific class. Let also ${\bm{X}}_{b}^{m}\in{\mathbb{R}}^{H\times W\times D}$ denote a baseline point in the uncertainty region of the model. We call it baseline because this point corresponds to an artificial image representation for which the prediction of the network is highly uncertain. Let ${l}_{b}^{m}\in{\mathbb{R}}$ denote the prediction logit for class $c$ that corresponds to ${\bm{X}}_{b}^{m}$, i.e. ${l}_{b}^{m}=f^{+}({\bm{X}}_{b}^{m})$. We define the following quantity:
${l}_{e}={l}_{c}-{l}_{b}^{m}$ (21)
[…]
5.2.2 Logit Difference Decomposition in terms of Concepts
Based on the aforementioned definitions, we have:
Due to the linearity of GAP, (23) can be written as:
Based on (27), the explanation logit ${l}_{e}$ is linearly expressed in terms of patches and concepts, giving detailed, spatially-aware information for concept contributions. We split (27) into four parts. The first part corresponds to contributions of concepts included in the sample (sample concept), the second to the contribution of concepts contained in an artificial baseline point of an uncertain prediction (baseline concept), the third part to a correction factor to account for imperfect direction learning convergence, and finally a residual that we do not explain. If the correction factor for concept $i$ was zero, concept $i$ would have a positive contribution to the explanation logit whenever the difference in contributions between the sample and the baseline is positive. In the case of imperfect learning conditions, this difference must exceed the negative of the correction factor.
In the experiments, we will be reporting Concept Contribution Maps (CCMs) where each spatial element $\bm{p}$} will be equal to:
visualized as a heatmap over the original image for a given concept $i$. Finally, we will be referring to the quantity ${\bm{w}}_{c}^{T}\hat{{\bm{s}}}_{i}$ as the Concept-Class Relation Coefficient (CCRC).
@@ 6 Experiment on Synthetic Data
Given the aforementioned synthetic dataset, we train a network to predict the image class based on the concept content of its patches. The network we use is composed of just two layers (corresponding to the top part of a potentially larger deep network). The first is an average-pooling layer, and the second is a linear layer with $K=3$ output classes. After training, the network attains 96.33% accuracy on a test set, randomly generated based on the previous principles. More details regarding the setup of this experiment are provided in Section A.4.
@@ 6.1 Evaluation of the Signal-Vector Estimator
Based on the synthetic pixel dataset $\{{\bm{x}}_{{\bm{p}}_{1}}\}\cup\{{\bm{x}}_{{\bm{p}}_{2}}\}$ and the ground-truth signal values ${\alpha}_{{\bm{p}},i}$ that we randomly generated, we put the estimator of Kindermans et al. (2017) and our signal vectors under test. The difference between the two estimators is the sub-sampling procedure that we proposed in Section 4.3. Given the ground-truth matrix of signal directions $\bm{S}$} we are able to calculate cosine similarities between the estimated signal directions and the respective ground-truth. The results of this experiment are depicted in Fig. 8 (in blue and green). Since we make use of the ground-truth signal values, this experiment evaluates the efficacy of signal vectors under perfect input conditions. The figure showcases the effectiveness of our proposed sub-sampling procedure when using (1) as in all cases signal vectors demonstrate perfect alignment with the ground-truth directions, while in the absence of sub-sampling, the estimation is less reliable. Conclusively, the sub-sampling technique that we proposed in Section 4.3 successfully adapts to the new assumptions that we make about the data.
Figure 8: 1) Cosine similarity between the ground-truth concept encoding directions and their estimation via Kindermans et al (1) calculated without our proposed feature sub-sampling. In this experiment, the estimation of signal directions uses ground-truth signal values. 2) Cosine similarity between the ground-truth concept encoding directions and their estimation via Signal-Vectors, i.e. using (1) calculated with our proposed feature sub-sampling. Similarly to the previous experiment this one uses ground-truth signal values. Under ideal input conditions, Signal-Vectors are proven to be faithful signal estimators. 3) Cosine similarity between ground-truth concept encoding directions and their estimation via our proposed Encoding-Decoding Direction Pairs (EDDP). In constrast to the previous two experiments, in this one Signal Vectors are learned in an unsupervised manner using all the aspects of the proposed method. Without access to ground-truth signal values, our overall unsupervised approach is capable of recovering the ground-truth concept encoding directions. 4) Cosine similarity between the ground-truth concept encoding directions and their estimation via our proposed Encoding-Decoding Direction Pairs (EDDP). This experiment is identical to (3) but omits the Filter-Signal Orthogonality Loss $\mathcal{L}^{fso}$ from the objective function. In that case, the encoding direction estimation is inferior compared to when employing $\mathcal{L}^{fso}$.

### v2

@@ 4.3 Signal Vectors as Concept Signal Estimators
In this paragraph, we consider the new data model that was outlined in Section 4.1 and try to validate whether (1) can be used to estimate the signal direction of a concept. Suppose that our objective is to estimate the signal direction of concept $i$ when we have access to a collection of patch embeddings $\{{\bm{x}}_{{\bm{p}}}\}$ and their signal values $\{\alpha_{{\bm{p}},i}\}$. Starting from the approach of PatternNet, it is easy to prove that whenever ${a}_{{\bm{p}},i}$ is independent of all ${a}_{{\bm{p}},j},j\neq i$ and ${\beta}_{{\bm{p}},f}$ the following property holds: $\text{cov}[{\bm{x}}_{{\bm{p}}}-{a}_{{\bm{p}},i}{\bm{s}}_{i},{a}_{{\bm{p}},i}]=0$, which, according to PatternNet directly implies that (1) can be utilized to estimate the direction of the signal.
@@ 4.4 Uncertainty Region Alignment to Discover Meaningful Concepts of Influence
To accomplish the alignment, all the patch embeddings ${\bm{x}}_{{\bm{p}}}$ in an image are manipulated towards the direction $\bm{dx_{{\bm{p}}}}$ to arrive at ${\bm{x}}_{{\bm{p}}}^{\prime}={\bm{x}}_{{\bm{p}}}-\bm{dx_{{\bm{p}}}}$. Based on our estimates of ${\bm{w}}_{i}$, $b_{i}$, and $\hat{{\bm{s}}}_{i}$, we select the direction $\bm{dx_{{\bm{p}}}}$ so that the shifted ${\bm{x}}_{{\bm{p}}}^{\prime}$ lies at the intersection of the concept detectors’ decision hyperplanes, i.e. ${\bm{w}}_{i}^{T}({\bm{x}}_{{\bm{p}}}-\bm{dx_{{\bm{p}}}})-b_{i}=0,\,\forall i$. Then, we ensure the network’s prediction for the resulting manipulated image representation is highly uncertain, effectively aligning both uncertainty regions. More specifically, we define two types of feature manipulation for this purpose:
Unconstrained Feature Manipulation (UFM) in which ${\bm{x}}_{{\bm{p}}}^{\prime}$ corresponds to the closest point of ${\bm{x}}_{{\bm{p}}}$ in the uncertainty region of the concept detectors, i.e.: ${\bm{x}}_{{\bm{p}}}^{\prime}=g_{{\bm{p}}}({\bm{x}}_{{\bm{p}}};{\bm{W}})={\bm{x}}_{{\bm{p}}}-\bm{dx_{{\bm{p}}}}$ and $\bm{dx_{{\bm{p}}}}$ given by:
$\bm{dx_{{\bm{p}}}}=({\bm{W}}^{T})^{+}({\bm{W}}^{T}{\bm{x}}_{{\bm{p}}}-{\bm{b}})$ (9)
with $\bm{A}$$^{+}$ denoting the pseudo-inverse of $\bm{A}$} and a detailed derivation provided in Section A.4.
Constrained Feature Manipulation (CFM) in which we additionally restrict feature manipulation to occur within the span of the signal vectors, i.e. $\bm{dx_{{\bm{p}}}}=\hat{{\bm{S}}}{\bm{v}}_{{\bm{p}}}$, ${\bm{v}}_{{\bm{p}}}\in{\mathbb{R}}^{I}$, with $\hat{{\bm{S}}}\in{\mathbb{R}}^{D\times I}$ denoting the matrix whose columns correspond to the learned signal vectors $\hat{{\bm{s}}}_{i},i\in\{0,1,...,I-1\}$. The feature manipulation formula for a patch embedding ${\bm{x}}_{{\bm{p}}}$ is $h_{{\bm{p}}}({\bm{x}}_{{\bm{p}}};{\bm{W}},\hat{{\bm{S}}})={\bm{x}}_{{\bm{p}}}-\bm{dx_{{\bm{p}}}}$, with ${\bm{v}}_{{\bm{p}}}$ and $\bm{dx_{{\bm{p}}}}$ given by:
${\bm{v}}_{{\bm{p}}}=({\bm{W}}^{T}\hat{{\bm{S}}})^{+}({\bm{W}}^{T}{\bm{x}}_{{\bm{p}}}-{\bm{b}})\Rightarrow$ (10)
@@
with $\mathcal{H}$} denoting entropy. When we manipulate the representations via either UFM or CFM, we use the layer’s activation function (typically a ReLU) to keep the features in the input domain of the next layer. Throughout the experiments, we will denote the Uncertainty Region Alignment loss with UFM as $\mathcal{L}^{uur}$ and with CFM as $\mathcal{L}^{cur}$. We consider the use of $\mathcal{L}^{uur}$ and $\mathcal{L}^{cur}$ to be mutually-exclusive. We either use the first or the second type of manipulation when computing the Uncertainty Region Alignment Loss and never both types at the same time.
Table 1: Encoding-Decoding Direction Pairs Loss Descriptions. Purple indicates loss contributions of this work, while light gray indicates loss terms from UIBE; CBE
Symbol Loss Name Loss Description
$\mathcal{L}^{fs}$ Focal Sparsity Ensure sparse patch-to-cluster assignments, emphasizing challenging cases
$\mathcal{L}^{ma}$ Maximum Activation Push towards binary patch-to-cluster memberships
$\mathcal{L}^{ic}$ Inactive Classifier Ensure all clusters have a sufficient number of members
$\mathcal{L}^{mm}$ Maximum Margin Ensure cluster members are linearly separated from non-cluster members by a sufficient margin
$\mathcal{L}^{eac}$ Excessively Active Classifier Penalize large clusters to avoid cases such as assigning all patches to a single cluster
$\mathcal{L}^{fso}$ Filter - Signal Orthogonality More accurate signal value extraction, enabling a more precise estimation of encoding directions
$\mathcal{L}^{ur}$ ($\mathcal{L}^{uur}$ or $\mathcal{L}^{cur}$) Uncertainty Region Alignment Discover interpretable concepts of influence exploiting network instructions.
@@ 4.5 Augmented Lagrangian Loss for Effective Direction Learning
The loss terms of the proposed method often conflict with each other. For instance, increasing confidence in the predictions of concept detectors ($\mathcal{L}^{ma}$) conflicts with maximizing the separation margin ($\mathcal{L}^{mm}$), and a large margin may also conflict with sparsity losses ($\mathcal{L}^{s}$ or $\mathcal{L}^{fs}$). Simply using a weighted sum of these losses requires careful architecture and dataset specific tuning of loss weights. The latter can be laborious and only indirectly controls clustering quality, which is mostly determined by the final loss values. To avoid this, we adopt an $\epsilon$-constrained optimization approach and reformulate our objective as a minimization problem with inequality constraints rather than linearly combining all loss terms. In this formulation, hyper-parameter tuning primarily pertains to target loss values, providing direct control over the clustering quality. We use the Augmented Lagrangian formulation (augmented-lagrangian1; augmented-lagrangian2) to solve the following optimization problem:
$\min\quad\lambda^{fs}\mathcal{L}^{fs}+\lambda^{ur}\mathcal{L}^{ur}\quad s.t.\quad\begin{cases}\mathcal{L}^{ma}\leq\tau^{ma}\\ \mathcal{L}^{ic}\leq\tau^{ic}\\ \mathcal{L}^{mm}\leq\tau^{mm}\\ \mathcal{L}^{eac}\leq\tau^{eac}\\ \mathcal{L}^{fso}\leq\tau^{fso}\end{cases}$ (13)
@@ 5.1 Reading Concept Information and Intervening on their Encoding
Reading Concept Information from Patch Embeddings: Reading concept information from a patch embedding is equivalent to estimating the signal value encoded in the representation. Filter directions can extract this value; however, this is done with a constant offset derived from the latent space bias $\bm{c}$}. This offset can be neutralized by calculating the difference between two projected embeddings (in the direction of the filter), which directly equals the difference in their signal values. Thus, reading concept information is helpful to be done with respect to the concept information in a point of reference. Two points of reference that we find useful are:
  • The Average Embedding: Estimating the signal value relative to a collection’s mean, for instance a collection of samples with or without the concept.
  • Point of Uncertainty: A point on the concept detector’s decision hyperplane that represents maximum concept ambiguity. Centering around this point ensures that positive values indicate concept presence and negative values indicate concept absence.
Intervening on Concept Encoding: Intervention modifies an embedding’s concept content by overwriting its signal value with a target value derived from a reference embedding. To encode the presence or absence of a concept in a patch embedding, we select target values based on:
  • The Average Signal Value in representative positive or negative concept samples.
  • Top or Bottom Quantiles of Signal Values in a collection of positive or negative concept samples.
Concrete implementation details for reading and intervening on concept information are provided in Section A.5.
@@
In this Section, we provide details on how the learned Encoding-Decoding Direction Pairs can be leveraged to provide concept-based local explanations for a network’s prediction. The core contribution is the introduction of an explanation logit (${l}_{e}={l}_{c}-{l}_{b}^{m}$), which is the difference between the network’s class logit for the input image (${l}_{c}$) and a class logit corresponding to a baseline artificial representation derived from the uncertainty region of the model (${l}_{b}^{m}$). We then decompose this ${l}_{e}$ into local, concept-specific components.
In Section A.6, we show that, for networks with a Global Average Pooling (GAP) penultimate layer, the explanation logit can be decomposed into four key components: the sample concept contributions, which are input dependent, the baseline concept contributions from an uncertain prediction, a correction factor compensating for imperfect learning convergence, and a final, unexplained residual.
Grouping the first two components in the previous decomposition and exploiting the linearity of GAP, the explanation logit enables the derivation of Concept Contribution Maps (CCMs), which provide a spatially-aware breakdown of the explanation logit in terms of its constituent concept contributions. CCMs differ from previous approaches (IBD; PCA) in that they do not simply localize the concepts in the input image, but they quantify the contributions of each concept across patches to the explanation logit, including both positive and negative contributions.
The contribution of a specific concept $i$ from an image patch $\bm{p}$} to the explanation logit is calculated as the product of two factors:
  • A Global Contribution Factor, referred to as Concept-Class Relation Coefficient (CCRC) (${\bm{w}}_{c}^{T}\hat{{\bm{s}}}_{i}$), independent of the individual image input, with ${\bm{w}}_{c}$ denoting the network’s class vector.
  • A Local Contribution Factor ($\hat{{v}}_{{\bm{p}},i}-\hat{{v}}_{{\bm{p}},i}^{b}$): A local, sample dependent, spatially-aware factor representing the difference in concept information between the actual image patch ($\hat{{v}}_{{\bm{p}},i}$) and the baseline patch ($\hat{{v}}_{{\bm{p}},i}^{b}$).
The CCM is a heatmap that visualizes the multiplication of these two factors across the image, precisely showing which regions and concepts positively or negatively drove the final classification decision. Further details are provided in Section A.6.
@@ 6 Experiment on Synthetic Data
Using this dataset, we train a network to predict the image class based on the concept content of its patches. The network we use is composed of just two layers (corresponding to the top part of a potentially larger deep network). The first is an average-pooling layer, and the second is a linear layer with $K=3$ output classes. After training, the network attains 96.33% accuracy on a test set, randomly generated based on the previous principles. More details regarding the setup of this experiment are provided in Section A.8.
@@ 6.1 Evaluation of the Signal-Vector Estimator
Based on the synthetic pixel dataset $\{{\bm{x}}_{{\bm{p}}_{1}}\}\cup\{{\bm{x}}_{{\bm{p}}_{2}}\}$ and the ground-truth signal values ${\alpha}_{{\bm{p}},i}$, we put the estimator of PatternNet and our signal vectors under test. The difference between the two estimators is the sub-sampling procedure that we proposed in Section 4.3. Given the ground-truth matrix of signal directions $\bm{S}$} we are able to calculate cosine similarities between the estimated signal directions and the respective ground-truth. The results of this experiment are depicted in Fig. 8 (in blue and green). Since we use the ground-truth signal values, this experiment evaluates the efficacy of signal vectors under perfect input conditions. The figure showcases the effectiveness of our proposed sub-sampling procedure when using (1), as in all cases, signal vectors demonstrate perfect alignment with the ground-truth directions, while in the absence of sub-sampling, the estimation is less reliable. Conclusively, the sub-sampling technique that we proposed in Section 4.3 adapts more precisely to our extended data model.
Figure 8: 1) Cosine similarity between the ground-truth concept encoding directions and their estimation via Kindermans et al (1) calculated without our proposed feature sub-sampling. In this experiment, the estimation of signal directions uses ground-truth signal values. 2) Cosine similarity between the ground-truth concept encoding directions and their estimation via Signal-Vectors, i.e. using (1) calculated with our proposed feature sub-sampling. Similarly to the previous experiment this one uses ground-truth signal values. Under ideal input conditions, Signal-Vectors are proven to be faithful signal estimators. 3) Cosine similarity between ground-truth concept encoding directions and their estimation via our proposed Encoding-Decoding Direction Pairs (EDDP). In constrast to the previous two experiments, in this one Signal Vectors are learned in an unsupervised manner using all the aspects of the proposed method. Without access to ground-truth signal values, our approach is capable of recovering the ground-truth concept encoding directions. 4) Cosine similarity between the ground-truth concept encoding directions and their estimation via our proposed Encoding-Decoding Direction Pairs (EDDP) but without using the Filter-Signal Orthogonality Loss $\mathcal{L}^{fso}$. In that case, the encoding direction estimation is inferior compared to (3).